If money is to quadruples in 13 years, what quarterly interest rate must the account earn?...

  1. If money is to quadruples in 13 years, what quarterly interest rate must the account earn? Assume quarterly compounding. Use formulas and equations to find the answer.
Answer #1

Let $ 1 is invested today ,so in 13 years Future value =$ 4

Number of quarters = 13 *4 = 52 quarters [4quarters in a year comprising of 3 months each]

using the formula ,

Future value =present value (1+i)^n

where i = quarterly interest rate

n=number of quarters

Now,putting the valurs in equation

4 =1 (1+i)^52

Multiplying both sides by ^1/52

(4)^1/52 = (1+i)^52*1/52

(4)^1/52 = (1+i)

(4)^.019231 = 1+i      [Find the value of (4)^.019231 using financial calculator )

1.027018 - 1 = =i

i = .027018 or 2.7018 %

quarterly interest rate = 2.70%